# Gradient Aligned Regression via Pairwise Losses

### v4

@@ 3.3 Efficient and Robust Aggregation for Different Losses
Given the huge diversity of the dataset and model, it is time-consuming to tune the three terms with a convex combination. Furthermore, the magnitude for each term can be different, which adds more difficulty to balance them. For example, $\mathcal{L}^{p=2}_{\text{diffnorm}}\in[0,2]$ but $\mathcal{L}^{\text{MAE}}_{c}$ can be infinitely large; therefore, $\mathcal{L}^{p=2}_{\text{diffnorm}}$ could be easily overwhelmed if simply combining other components with larger magnitudes by arithmetic mean, i.e. $\mathcal{L}=(\mathcal{L}^{\text{MAE}}_{c}+\mathcal{L}^{\text{MSE}}_{\text{diff}}+\mathcal{L}^{p=2}_{\text{diffnorm}})/3$.
One intuitive way for amplifying the significance of $\mathcal{L}^{p=2}_{\text{diffnorm}}$ is geometric mean, i.e. $\mathcal{L}=(\mathcal{L}^{\text{MAE}}_{c}\mathcal{L}^{\text{MSE}}_{\text{diff}}\mathcal{L}^{p=2}_{\text{diffnorm}})^{1/3}$, where the effects of different loss magnitudes are unified. However, it could bring another issue. Consider $\mathcal{L}(a,b,c)=(abc)^{1/3}$, $\frac{\partial\mathcal{L}(a,b,c)}{\partial a}=\frac{(bc)^{1/3}}{a^{2/3}}$. The smaller component can get larger gradient but may get numerical issue when $a\ll bc$. As a consequence, the overall objective could solely focuses on the smallest component.
Inspired by the previous two intuitive examples (the loss with smaller magnitude either gains too less attention or too much attention), we propose the an efficient and robust aggregation approach for GAR based on a variant of Distributionally Robust Optimization (DRO), which can not only trade off the previous two cases but also only requires one tuning hyper-parameter (Zhu et al., 2023). For each loss, we apply a logarithmic transformation for reducing magnitude effect. For the sake of generality, we denote the loss as $\mathcal{L}_{i}$ and the overall loss as $\mathcal{L}_{\text{GAR}}(\mathcal{L}_{1},...,\mathcal{L}_{M})$, where the overall loss is composed with $M$ sub-losses. Denote $D(\mathbf{p}|\frac{\mathbf{1}}{M})$ as divergence measure from probability vector $\mathbf{p}$ over simplex $\Delta_{M}$ to the uniform distribution $\frac{\mathbf{1}}{M}$. The DRO formulation for GAR takes the balance from the averaged value to the maximal value:
@@
$\exp\big{(}\mathcal{L}^{\text{KL}}_{\text{GAR}}$ $(\mathcal{L}_{1},...,\mathcal{L}_{M};\alpha)\big{)}=(\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}_{i}^{1/\alpha})^{\alpha}=\begin{cases}(\Pi_{i=1}^{M}\mathcal{L}_{i})^{\frac{1}{M}}~{},&\alpha\rightarrow+\infty.\\ \frac{1}{M}\sum_{i=1}^{M}\mathcal{L}_{i}~{},&\alpha=1.\\ \arg\max_{i=1}^{M}\mathcal{L}_{i}~{},&\alpha\rightarrow 0.\\ \end{cases}$ (10)
@@
Next, we discuss how to take care of extreme case for $\mathcal{L}_{\text{GAR}}$. Notice that $\mathcal{L}_{i}^{1/\alpha}$ can have numerical issue when $\alpha\rightarrow 0,\mathcal{L}_{i}\rightarrow+\infty$ or $\alpha\rightarrow+\infty,\mathcal{L}_{i}\rightarrow 0$. The former case can cause computation overflow on forward propagation and the later case can cause computation overflow on backward propagation. Denote $\mathcal{L}_{\text{max}}=\arg\max_{i=1}^{M}\mathcal{L}_{i}$ and $\mathcal{L}_{\text{min}}=\arg\min_{i=1}^{M}\mathcal{L}_{i}$. We utilize the constant value (detached from back propagation) of $L_{\text{max}}$ or $L_{\text{min}}$ to control $L^{\text{KL}}_{\text{GAR}}$. It is easy to show $\mathcal{L}^{\text{KL}}_{\text{GAR}}$ has the following equivalent empirical formulation:
@@ 4 Experiments
We demonstrate the effectiveness of GAR with two synthetic datasets, eight real-world tasks on five tabular benchmark datasets and one image benchmark datasets. Further analysis for ablation studies on different components, hyper-parameter $\alpha$ and running time of GAR are presented afterwards.
@@ 4.1 Synthetic Experiments
where recall that $\mathcal{L}^{p=2}_{\text{diffnorm}}=1-\rho(f,y)$, and $\beta$ is dynamically assigned as the constant value of $\rho(f,y)$ on each iteration. Intuitively, when the Pearson correlation coefficient is high, the model focuses more on the MAE loss; and vice versa. It is worth noting that GAR can be also adapted to Symbolic Regression (SR) settings. For example, GAR loss could be defined as fitness function for evolutionary algorithm based method or reward function for reinforcement learning based method  (Cranmer, 2023; Petersen et al., 2019). However, given that this work focuses on Numeric Regression (NR) setting, it is unfair to directly compare the proposed GAR under NR setting with SR methods here. Because SR methods have put the ground truth functions in their ‘toolbox’ (as search candidates) for these toy examples. To elaborate this point, we extensively employ a well-established SR method, PySR [^1] https://astroautomata.com/PySR/ on the synthetic datasets. We find PySR performs perfectly when trigonometric functions are included in the search space. However, it performs much worse when we exclude trigonometric functions. Results are presented in the Appendix B.5.
@@
Experimental Settings: For both of the synthetic datasets, we uniformly randomly sample half of the data as the training dataset. The full details for experimental settings are in the Appendix B.3.
@@ 4.3 Further Analysis
Table 2: Running time (seconds/epoch) on five benchmark datasets. Repeated 1500 epochs, mean and standard deviation are reported. ‘CCS’: Concrete Compressive Strength. ’All Time’: all the time consumed per epoch. ’Loss Time’: the time for loss calculation and loss backward operation.
@@ Acknowledgments
A Proofs
@@ A.1 Proof for Theorem 1
Recall that $\bar{\sigma}_{\mathbf{x}}^{f}=\frac{1}{N}\sum_{i=1}^{N}\sigma_{\mathbf{x}_{i}}^{f}$ denotes the empirical mean.
@@
$=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}(\sigma_{\mathbf{x}_{i}}^{f}-\sigma_{\mathbf{x}_{j}}^{f})^{2}$
$=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}(\sigma_{\mathbf{x}_{i}}^{f}-\bar{\sigma}_{\mathbf{x}}^{f}+\bar{\sigma}_{\mathbf{x}}^{f}-\sigma_{\mathbf{x}_{j}}^{f})^{2}$
$=\frac{1}{N}\sum_{i=1}^{N}(\sigma_{\mathbf{x}_{i}}^{f}-\bar{\sigma}_{\mathbf{x}}^{f})^{2}$
$=\text{Var}(\sigma_{\mathbf{x}}^{f}),$
@@ B.2 Data Pre-processing for IC50
B.3 Experimental Settings on Synthetic Datasets
For both of the synthetic datasets, we uniformly randomly sample half of the data as the training dataset. We utilize a 7-layer simple Feed Forward Neural Network (FFNN) with 5 hidden layers where neuron number for each layer is set as 100. We utilize ELU function as the activation function (Clevert et al., 2015). The total training epochs are set as 300 and batch size is set as 128. The initial learning rate is tuned in {1e-1, 1e-2, 1e-3, 1e-4} for Adam optimizer (Kingma and Ba, 2014), which is stage-wised decreasing by 10 folds at the end of the 100-th and 200-th epoch. The weight decay is tuned in {1e-3,1e-4,1e-5,0}. RNC takes the first 100 epochs are pre-training and the remaining as fine-tuning with MAE loss; temperature for RNC is tuned in {1,2,4}. For each baseline, we run 5 trials independently with the following 5 seeds: {1,2,3,4,5}. We set the $\alpha=0.5$ for GAR on all synthetic experiments.
B.4 MSE on Synthetic Datasets
@@
B.5 Symbolic Regression on the Synthetic Datasets
@@
B.6 Backbone Model Complexity Variation on Sine Dataset
@@
B.7 Experimental Setting on Real World Datasets
@@
Table 4: The selected $\alpha$ on the tabular data. Standard deviations are included in the parenthesis.
@@
B.8 Experiment Settings for AgeDB
We follow the the same experiment settings from the previous work (Zha et al., 2023). The age range is between 0 and 101. It is split into 12208/2140/2140 images for training/validation/testing sets. The SGD optimizer and cosine learning rate annealing is utilized for training (Loshchilov and Hutter, 2016). The batch size is set to 256. For both training from scratch and linear probe experiments, we select the best learning rates and weight decays for each dataset by grid search, with a grid of learning rates from {0.01, 0.05, 0.1, 0.2, 0.5, 1.0} and weight decays from {1e-6, 1e-5, 1e-4, 1e-3}. For the predictor training of two-stage methods, we adopt the same search setting as above except for adding additional 0 weight decay to the search choices of weight decays. The RNC temperature parameter is search from {0.1, 0.2, 0.5, 1.0, 2.0, 5.0} and select the best, which is 2.0. We train all one-stage methods and the encoder of two-stage methods for 400 epochs, and the linear regressor of two-stage methods for 100 epochs. The hyper-parameters for all the compared methods are searched in the same region as described in appendix B.7. For training from scratch experiments, because the ConR (Keramati et al., 2023) is proposed to use two data augmentations, we utilize two data augmentations for each training data for all compared methods. The $\alpha$ for GAR is selected as 0.1 for training from scratch setting and as 1.0 for linear probe setting. For each method, we randomly repeat five times with different random seeds to report the mean and standard deviation for evaluations.
B.9 Ablation Studies on GAR
@@ B.10 Sensitivity Analysis
B.11 Different Batch Sizes
@@
B.12 Difference with Standardization

### v6

@@ 3.3 Efficient and Robust Aggregation for Different Losses
Given the huge diversity of the dataset and model, it is time-consuming to tune the three terms with a convex combination. Furthermore, the magnitudes for each term can be different, which adds more difficulty to balance them. For example, $\mathcal{L}^{p=2}_{\text{diffnorm}}\in[0,2]$ but $\mathcal{L}^{\text{MAE}}_{c}$ can be infinitely large; therefore, $\mathcal{L}^{p=2}_{\text{diffnorm}}$ could be easily overwhelmed if simply combining other components with larger magnitudes by arithmetic mean, i.e. $\mathcal{L}=(\mathcal{L}^{\text{MAE}}_{c}+\mathcal{L}^{\text{MSE}}_{\text{diff}}+\mathcal{L}^{p=2}_{\text{diffnorm}})/3$.
An intuitive way to amplify the significance of $\mathcal{L}^{p=2}_{\text{diffnorm}}$ is the geometric mean, that is, $\mathcal{L}=(\mathcal{L}^{\text{MAE}}_{c}\mathcal{L}^{\text{MSE}}_{\text{diff}}\mathcal{L}^{p=2}_{\text{diffnorm}})^{1/3}$, where the effects of different loss magnitudes are unified. However, it could raise another issue. Consider $\mathcal{L}(a,b,c)=(abc)^{1/3}$, $\frac{\partial\mathcal{L}(a,b,c)}{\partial a}=\frac{(bc)^{1/3}}{a^{2/3}}$. The smaller component can get a larger gradient but may experience a numerical issue when $a\ll bc$. As a consequence, the overall objective could focus solely on the smallest component.
Inspired by the previous two intuitive examples (the loss with smaller magnitude either gains too less attention or too much attention), we propose an efficient and robust aggregation approach for GAR based on a variant of Distributionally Robust Optimization (DRO), which can not only trade off the previous two cases, but also only requires one tuning hyperparameter (Zhu et al., 2023). For each loss, we apply a logarithmic transformation to reduce the magnitude effect. For the sake of generality, we denote the loss as $\mathcal{L}_{i}$ and the overall loss as $\mathcal{L}_{\text{GAR}}(\mathcal{L}_{1},...,\mathcal{L}_{M})$, where the overall loss is composed with $M$ sub-losses. Denote $D(\mathbf{p}|\frac{\mathbf{1}}{M})$ as divergence measure from probability vector $\mathbf{p}$ over simplex $\Delta_{M}$ to the uniform distribution $\frac{\mathbf{1}}{M}$. The DRO formulation for GAR takes the balance from the averaged value to the maximal value:
@@
$\exp\big{(}\mathcal{L}^{\text{KL}}_{\text{GAR}}$ $(\mathcal{L}_{1},...,\mathcal{L}_{M};\alpha)\big{)}=(\frac{1}{M}\sum_{i=1}^{M}\mathcal{L}_{i}^{1/\alpha})^{\alpha}$
=$=\begin{cases}(\Pi_{i=1}^{M}\mathcal{L}_{i})^{\frac{1}{M}}~{},&\alpha\rightarrow+\infty.\\ \frac{1}{M}\sum_{i=1}^{M}\mathcal{L}_{i}~{},&\alpha=1.\\ \arg\max_{i=1}^{M}\mathcal{L}_{i}~{},&\alpha\rightarrow 0.\\ \end{cases}$ (10)
@@
Next, we discuss how to take care of extreme case for $\mathcal{L}_{\text{GAR}}$. Notice that $\mathcal{L}_{i}^{1/\alpha}$ can have numerical issue when $\alpha\rightarrow 0,\mathcal{L}_{i}\rightarrow+\infty$ or $\alpha\rightarrow+\infty,\mathcal{L}_{i}\rightarrow 0$. The former case can cause computation overflow on forward propagation and the later case can cause computation overflow on backward propagation. Denote $\mathcal{L}_{\text{max}}=\arg\max_{i=1}^{M}\mathcal{L}_{i}$ and $\mathcal{L}_{\text{min}}=\arg\min_{i=1}^{M}\mathcal{L}_{i}$. We utilize the constant value (detached from back propagation) of $\mathcal{L}_{\text{max}}$ or $\mathcal{L}_{\text{min}}$ to control $\mathcal{L}^{\text{KL}}_{\text{GAR}}$. It is easy to show $\mathcal{L}^{\text{KL}}_{\text{GAR}}$ has the following equivalent empirical formulation:
@@ 4 Experiments
We demonstrate the effectiveness of GAR with two synthetic datasets, eight real-world tasks on five tabular benchmark datasets and one image benchmark dataset. Further analysis for running time of GAR, ablation studies on different components, sensitivity to hyper-parameter $\alpha$ and batch sizes are presented afterwards.
@@ 4.1 Synthetic Experiments
where recall that $\mathcal{L}^{p=2}_{\text{diffnorm}}=1-\rho(f,y)$, and $\beta$ is dynamically assigned as the constant value of $\rho(f,y)$ on each iteration. Intuitively, when the Pearson correlation coefficient is high, the model focuses more on the MAE loss; and vice versa. It is worth noting that GAR can also be adapted to Symbolic Regression (SR) settings. For example, GAR loss could be defined as a fitness function for evolutionary algorithm based method or reward function for reinforcement learning based method  (Cranmer, 2023; Petersen et al., 2019). However, given that this work focuses on Numeric Regression (NR) setting, it is unfair to directly compare the proposed GAR under the NR setting with SR methods here. Because SR methods have placed the ground truth functions in their ‘toolbox’ (as search candidates) for these toy examples. To elaborate on this point, we extensively employ a well-established SR method, PySR [^1] https://astroautomata.com/PySR/ on the synthetic datasets. We find PySR performs perfectly when trigonometric functions are included in the search space. However, it performs much worse when we exclude trigonometric functions. Results are presented in the Appendix B.4.
@@
Experimental Settings: For both of the synthetic datasets, we uniformly randomly sample half of the data as the training dataset. The total training epochs are set as 300 and batch size is set as 128. The initial learning rate is tuned in {1e-1, 1e-2, 1e-3, 1e-4} for Adam optimizer (Kingma & Ba, 2014), which is stage-wised decreasing by 10 folds at the end of the 100-th and 200-th epoch. The weight decay is tuned in {1e-3,1e-4,1e-5,0}. RNC takes the first 100 epochs are pre-training and the remaining as fine-tuning with MAE loss; temperature for RNC is tuned in {1,2,4}. For each baseline, we run 5 trials independently with the following 5 seeds: {1,2,3,4,5}. We set the $\alpha=0.5$ for GAR on all synthetic experiments.
@@ 4.3 Further Analysis
Table 2: Running time (seconds/epoch) on five benchmark datasets. Repeated 1500 epochs, mean and standard deviation are reported. ‘CCS’: Concrete Compressive Strength. ‘All Time’: all the time consumed per epoch. ‘Loss Time’: the time for loss calculation and loss backward operation.
@@ Acknowledgement
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proofs
@@ A.1 Proof for Theorem 1
Recall that $\bar{\delta}_{\mathbf{x}}^{f}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{x}_{i}}^{f}$ denotes the empirical mean.
@@
$=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}(\delta_{\mathbf{x}_{i}}^{f}-\delta_{\mathbf{x}_{j}}^{f})^{2}$
$=\frac{1}{2N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}(\delta_{\mathbf{x}_{i}}^{f}-\bar{\delta}_{\mathbf{x}}^{f}+\bar{\delta}_{\mathbf{x}}^{f}-\delta_{\mathbf{x}_{j}}^{f})^{2}$
$=\frac{1}{N}\sum_{i=1}^{N}(\delta_{\mathbf{x}_{i}}^{f}-\bar{\delta}_{\mathbf{x}}^{f})^{2}$
$=\text{Var}(\delta_{\mathbf{x}}^{f}),$
@@ B.2 Data Pre-processing for IC50
B.3 MSE on Synthetic Datasets
@@
B.4 Symbolic Regression on the Synthetic Datasets
@@
B.5 Backbone Model Complexity Variation on Sine Dataset
@@
B.6 Experimental Setting on Real World Datasets
@@
Table 4: The selected $\alpha$ on the tabular data respecting to different evaluation metrics. Standard deviations are included in the parenthesis.
@@
B.7 Experiment Settings for AgeDB
We follow the the same experiment settings from the previous work (Zha et al., 2023). The age range is between 0 and 101. It is split into 12208/2140/2140 images for training/validation/testing sets. The SGD optimizer and cosine learning rate annealing is utilized for training (Loshchilov & Hutter, 2016). The batch size is set to 256. For both training from scratch and linear probe experiments, we select the best learning rates and weight decays for each dataset by grid search, with a grid of learning rates from {0.01, 0.05, 0.1, 0.2, 0.5, 1.0} and weight decays from {1e-6, 1e-5, 1e-4, 1e-3}. For the predictor training of two-stage methods, we adopt the same search setting as above except for adding additional 0 weight decay to the search choices of weight decays. The RNC temperature parameter is search from {0.1, 0.2, 0.5, 1.0, 2.0, 5.0} and select the best, which is 2.0. We train all one-stage methods and the encoder of two-stage methods for 400 epochs, and the linear regressor of two-stage methods for 100 epochs. The hyper-parameters for all the compared methods are searched in the same region as described in appendix B.6. For training from scratch experiments, because the ConR (Keramati et al., 2023) is proposed to use two data augmentations, we utilize two data augmentations for each training data for all compared methods. The $\alpha$ for GAR is selected as 0.1 for training from scratch setting and as 1.0 for linear probe setting. For each method, we randomly repeat five times with different random seeds to report the mean and standard deviation for evaluations.
B.8 Ablation Studies on GAR
@@ B.9 Sensitivity Analysis
B.10 Different Batch Sizes
@@
B.11 Difference with Standardization